Omar Industries manufactures two products: Regular and Super. The results of operations for 20x1 follow. Regular Super Total Units 12,000 3,900 15,900 Sales revenue $ 312,000 $ 819,000 $ 1,131,000 Less: Cost of goods sold 240,000 468,000 708,000 Gross Margin $ 72,000 $ 351,000 $ 423,000 Less: Selling expenses 72,000 180,000 252,000 Operating income (loss) $ 0 $ 171,000 $ 171,000 Fixed manufacturing costs included in cost of goods sold amount to $4 per unit for Regular and $20 per unit for Super. Variable selling expenses are $5 per unit for Regular and $20 per unit for Super; remaining selling amounts are fixed. Omar Industries wants to drop the Regular product line. If the line is dropped, company-wide fixed manufacturing costs would fall by 10% because there is no alternative use of the facilities. What would be the impact on operating income if Regular is discontinued?
Answer #1

Total fixed manufacturing cost = 4*12,000 + 20*3,900 = $126,000

Impact on operating income if regular is discontinued = savings in fixed costs - contribution margin lost

= 10%*126,000 - [312,000- (240,000-4*12,000) - 12,000*5]

= 12,600 - 60,000

= - $47,400

Hence, the operating income will fall by $47,400